# InPK: Infusing Prior Knowledge into Prompt for Vision-Language Models

### v1

@@ 3.2.4 Training Objective
where $\operatorname{sim}(\cdot,\cdot)$ denotes the cosine similarity, $\tau$ is a temperature parameter that controls the sharpness of the softmax distribution, and $N_{c}$ the number of seen classes. Thus, the cross-entropy loss between the true label and the predicted label is minimized as follows:
@@
Furthermore, the attribute words of different categories are often generic. For example, Felidae species such as the cat and tiger share common fine-grained attributes, including the presence of whiskers, a long tail, and so on. The constraints of $\mathcal{L}_{ce}$ alone would cause the prompt to become largely dependent on the attribute words, potentially confusing the categories. It is therefore necessary to enhance the role of class names by leveraging the differences of class names within the general text features $\boldsymbol{f}$. Following the PromptSRC [14] approach, we introduce a regularization constraint using the $L1$ Loss on the prompted textual features $\boldsymbol{\widetilde{f}}$ to ensure their consistency with the CLIP pre-trained features $\boldsymbol{f}$. This regularization prevents the model from overfitting to overly generic or misleading attributes, ensuring that the generated features align with the pre-trained model’s feature space, without consuming additional inference time. The loss is minimized as follows:
@@
$\mathcal{L}=\mathcal{L}_{ce}+\lambda\mathcal{L}_{text},$ (13)
where $\lambda$ is a hyperparameter set to 25.0.
4 Experiments
4.1 Experimental Setup
We follow the experimental setting [50] of previous methods to conduct experiments through three evaluation protocols: base-to-novel generalization, few-shot learning, and cross-dataset evaluation.
@@ 4.1 Experimental Setup
(a) Average over 11 datasets.
@@
(b) ImageNet
@@
CoOp 93.67 95.29 94.47
CoCoOp 95.20 97.69 96.43
TCP 94.67 97.20 95.92
Ours 94.67 97.63 96.13
$\Delta$ +0.00 +0.43 +0.21
(d) OxfordPets
Method Base Novel HM
CoOp 78.12 60.40 68.13
CoCoOp 70.49 73.59 72.01
TCP 80.80 74.13 77.32
Ours 81.97 76.10 78.93
$\Delta$ +1.17 +1.97 +1.60
(e) StanfordCars
@@
(f) Flowers102
Method Base Novel HM
CoOp 88.33 82.26 85.19
CoCoOp 90.70 91.29 90.99
TCP 90.57 91.37 90.97
Ours 90.10 91.57 90.83
$\Delta$0.47 +0.20 -0.14
(g) Food101
Method Base Novel HM
CoOp 40.44 22.30 28.75
CoCoOp 33.41 23.71 27.74
TCP 41.97 34.43 37.83
Ours 45.67 36.50 40.57
$\Delta$ +3.70 +2.07 +2.75
(h) FGVCAircraft
@@
(i) SUN397
Method Base Novel HM
CoOp 79.44 41.18 54.24
CoCoOp 77.01 56.00 64.85
TCP 82.77 58.07 68.25
Ours 84.43 71.30 77.31
$\Delta$ +1.66 +13.23 +9.06
(j) DTD
Method Base Novel HM
CoOp 92.19 54.74 68.69
CoCoOp 87.49 60.04 71.21
TCP 91.63 74.73 82.32
Ours 95.70 82.37 88.54
$\Delta$ +4.07 +7.64 +6.21
(k) EuroSAT
Method Base Novel HM
CoOp 84.69 56.05 67.46
CoCoOp 82.33 73.45 77.67
TCP 87.13 80.77 83.83
Ours 87.13 83.10 85.07
$\Delta$ +0.00 +2.33 +1.24
(l) UCF101
Table 1: Comparison with CoOp, CoCoOp and TCP in the base-to-novel generalization setting. HM: harmonic mean. The best results for each column are highlighted in bold font, and the improvement of our method over the SOTA method TCP is indicated as $\Delta$. Our method exhibits outstanding performance in both base and novel classes.
CoOp [50] CoCoOp [49] PromptSRC [14] MaPLe [13] MMA [36] TCP  [38] HPT [28] ProMetaR [23] Ours
(IJCV22) (CVPR22) (ICCV23) (CVPR23) (CVPR24) (CVPR24) (AAAI24) (CVPR24)
Base 82.69 80.47 84.26 82.28 83.20 84.13 84.32 84.39 85.17
Novel 63.22 71.69 76.10 75.14 76.80 75.36 76.86 76.93 78.79
HM 71.66 75.83 79.97 78.55 79.87 79.51 80.23 80.49 81.85
Table 2: Comparison between our method and the existing state-of-the-art methods in the base-to-novel generalization setting. Our method achieves the best accuracy in both base and novel classes, demonstrating the strong generalization ability of our method.
Method ImNet Caltech Pets Cars Flowers Food Aircraft SUN397 DTD EuroSAT UCF Average
CoOp 66.70 93.30 89.40 65.70 70.70 85.90 24.90 62.60 44.30 48.30 67.60 65.37
CoCoOp 70.55 94.98 93.01 69.10 82.56 86.64 30.87 70.50 54.79 63.83 74.99 71.98
PromptSRC 70.80 94.77 93.23 71.83 91.31 86.06 32.80 72.80 60.64 75.02 79.35 75.33
MaPLe 70.67 94.30 92.05 68.70 80.80 86.90 29.03 71.47 54.73 54.87 73.70 70.66
TCP 70.48 95.00 91.90 76.30 94.40 85.30 36.20 72.11 63.97 77.43 80.83 76.72
Ours 71.00 95.00 92.90 77.57 93.27 86.10 36.23 73.73 64.77 83.20 80.30 77.64
Table 3: Comparison of our method with existing methods in the few-shot learning setting. All models are trained with 4-shot samples. Our method shows competitive performance in few-shot learning, achieving the highest average results.
@@
Table 4: Comparison of InPK and existing methods in cross-data evaluation settings. All models are trained with 16-shot samples on the source dataset ImageNet and evaluated on 10 other datasets. Our method achieves the highest average accuracy, indicating that InPK has excellent robustness to different data distributions.
@@
Few-shot classification. To evaluate the model’s ability to learn with extremely limited supervision, we assess its performance across few-shot scenarios. Specifically, the model is trained on a small number of labeled images (4-shot) and evaluated on a dataset containing the same classes as those in the training samples.
@@
Experiment detail. Our implementation is based on the code of CoOp [50]. All experiments are fine-tuned using the CLIP model with a ViT-B/16 backbone where $d=512$, and the results are averaged over three runs with different random seeds (1/2/3). For more experimental details, please refer to the supplementary material.
4.2 Base-to-Novel Generalization
@@
Tab. 1 presents the performance of InPK across 11 recognition datasets in the base-to-novel generalization setting. We implement all methods using a few-shot training approach, with 16 randomly sampled shots per base class. Compared to the recently proposed method TCP [38], our approach shows enhanced performance on 10 out of 11 datasets in terms of harmonic mean (HM), with only a slight drop in performance on the Food101 dataset relative to method TCP. For the average result, since we infuse class-specific prior knowledge into learnable tokens in Pki, these tokens are expected to learn the common features shared across modalities. This allows our method to perform better on novel class, improving the average accuracy of novel class by 3.43%. Additionally, InPK maintains strong performance on the base class as it can infer more discriminative text features, with an increase of 1.04% in accuracy. When considering both base and novel classes, InPK achieves an absolute average gain of 2.34% over TCP in terms of HM, demonstrating an effective balance between in-domain and out-of-domain data performance.
@@
Tab. 2 presents a comparison of the average performance of our method against more recent prompt-tuning methods across 11 datasets. The results evident that our method outperforms other methods on both base and novel classes, which demonstrates that the InPK not only maintains excellent recognition accuracy on base class but also exhibits strong generalization capabilities for novel class.
4.3 Few-Shot Learning
Tab. 3 shows the performance of InPK in few-shot learning across 11 datasets. Our method achieves the highest gains on most datasets (7 out of 11). Compared to the best existing method TCP, our model achieves an average improvement of 0.92%. Since InPK introduces class prior knowledge, it compensates for the small number of samples by utilizing this prior knowledge as additional useful information, enhancing the model’s perception of fine-grained differences. This demonstrates that InPK maintains high performance even in limited data scenarios.
4.4 Cross-Dataset Evaluation
@@
To further verify the generalization ability of the model, we evaluate the model and state-of-the-art methods in a cross-data setting. As shown in Tab. 4, our model exhibits outstanding generalization on 8 out of the 10 datasets, with a significant improvement over other methods. The average accuracy reaches 69.90%, demonstrating strong robustness across varying data distributions. This is because the model infuses the prior knowledge into learnable tokens at various feature levels, which guides the model to focus on category information and better discover latent representations of novel class attributes, thereby improving the model’s performance on unseen classes.
4.5 Ablation Study
@@
Impact of prompt length. We study the effect of prompt length $M$ on model performance in a base-to-novel generalization setting. As shown in Fig. 3 (left), changing $M$ has minimal impact on base class performance. However, for the novel class, performance improves as the prompt length increases. Ultimately, we select $M=6$, which yields the highest harmonic mean (HM) accuracy, as the optimal prompt length for the model.
@@
Impact of the number of attribute words. We analyze the impact of the number of attribute words $N$ on model performance in the base-to-novel generalization setting as shown in Fig. 3 (right). By varying the number of attribute words, we observe the accuracy of the base class, new class, and harmonic mean (HM). Overall, as the number of attribute words increases, the HM index gradually improves. This is because richer class-specific textual knowledge provides more comprehensive supervision for the model. Notably, when the number of attribute words $N=25$, the HM value is the highest, indicating that this amount of detailed prior knowledge optimizes the balance between base and novel class performance.
@@
Impact of inserting attribute-aware attention into different Transformer layers. Tab. 5 presents the analysis of inserting attribute-aware attention into different Transformer layers within PKi in the base-to-new generalization setting. The results indicate that even with $J=1$, where attribute-aware attention is only applied in the initialization stage, the model’s performance shows a significant improvement. This highlights the importance of providing a prior knowledge-based initialization for the learnable tokens. The model performance is observed to improve with the increase of the prompt depth of module insertion, which can be attributed to the fact that class-specific prior knowledge is able to reduce the loss of category information. Additionally, inserting attribute-aware attention beyond the ninth layer decreases performance, likely because earlier layers have already captured high-level features. Further insertions may interfere with these representations, making the model overly sensitive and reducing its generalization ability. The optimal result is obtained by incorporating attribute-aware attention into the transformer layers up to a depth $J=9$.
@@
Table 5: Impact of insert layers.
@@
Analysis of model components. We conduct a comprehensive analysis of the key components of InPK, as illustrated in Tab. 6. In this analysis, Know. represents the input attribute words, Attn. refers to the attribute-aware attention, and Proj. represents the learnable projection layer from text prompt to visual prompt. In the text branch, when only class-specific prior knowledge is applied, the model exhibits improvement with an average gain from 79.03% to 80.69% due to the additional supervision provided by the rich attribute information. On the other hand, when only attribute-aware attention is used, the model strikes a better balance between base and novel classes with an average improvement of 0.92%. When both components (Know. and Attn.) are incorporated together, the model achieves substantial gains across both base and new classes (83.70% vs 85.17%, 74.86% vs 78.79%). In the visual branch, incorporating the projection layer improves the HM accuracy by 1.05% compared to the setup without the projection layer. This gain is primarily due to the projection layer’s ability to enhance consistency between the image and text semantics, further optimizing cross-modal alignment.
Text encoder Image encoder Base Novel HM
Know. Attn. Proj.
@@
Table 6: Ablation on attribute words, attribute-aware attention and learnable text-to-vision projection layer in InPK.
Effectiveness of prior knowledge infusion. To further assess the impact of prior knowledge infusion (pki) into learnable tokens, we apply pki only before the first encoder layer in method CoOp [50], KgCoOp [37], and PromptSRC [14]. As shown in Tab. 7, all three methods exhibit significant improvement, confirming the effectiveness of pki.
@@
+ pki 85.15+0.89 78.60+2.50 81.74+1.77
Table 7: Comparison of baseline methods with and without pki in the base-to-novel generalization setting. ‘pki’ denotes prior knowledge infusion into learnable tokens before feeding them into the encoder layer.
5 Conclusion

### v2

@@ III-B4 Training Objective
where $\operatorname{sim}(\cdot,\cdot)$ denotes the cosine similarity, $\tau$ is a temperature parameter that controls the sharpness of the softmax distribution, and $N_{c}$ is the number of seen classes. Thus, the cross-entropy loss between the true label and the predicted label is minimized as follows:
@@
Furthermore, the attribute words of different categories are often generic. For example, Felidae species such as the cat and tiger share common fine-grained attributes, including the presence of whiskers, a long tail, and so on. The constraints of $\mathcal{L}_{ce}$ alone would cause the prompt to become largely dependent on the attribute words, potentially confusing the categories. It is therefore necessary to enhance the role of class names by leveraging the differences of class names within the general text features $\boldsymbol{f}$. Following the PromptSRC [7] approach, we introduce a Class-Name Saliency Regularization constraint using the $L1$ Loss on the prompted textual features $\boldsymbol{\widetilde{f}}$ to ensure their consistency with the CLIP pre-trained features $\boldsymbol{f}$. This regularization prevents the model from overfitting to overly generic or misleading attributes, ensuring that the generated features align with the pre-trained model’s feature space without consuming additional inference time. The loss is minimized as follows:
@@
$\mathcal{L}=\mathcal{L}_{ce}+\lambda\mathcal{L}_{text}.$ (14)
IV Experiments
IV-A Experimental Setup
We follow the experimental setting [4] of previous methods to conduct experiments through four evaluation protocols: base-to-novel generalization, few-shot classification, cross-dataset evaluation, and domain generalization.
TABLE I: Comparison with CoOp, CoCoOp and TCP in the base-to-novel generalization setting. HM: harmonic mean. The best results for each column are highlighted in bold font, and the improvement of our method over the SOTA method TCP is indicated as $\Delta$. Our method exhibits outstanding performance in both base and novel classes.
(a) Average results
@@ IV-A Experimental Setup
(d) DTD
Method Base Novel HM
CoOp 79.44 41.18 54.24
CoCoOp 77.01 56.00 64.85
TCP 82.77 58.07 68.25
Ours 84.43 71.30 77.31
$\Delta$ +1.66 +13.23 +9.06
(g) OxfordPets
Method Base Novel HM
CoOp 93.67 95.29 94.47
CoCoOp 95.20 97.69 96.43
TCP 94.67 97.20 95.92
Ours 94.67 97.63 96.13
$\Delta$ +0.00 +0.43 +0.21
(j) Food101
Method Base Novel HM
CoOp 88.33 82.26 85.19
CoCoOp 90.70 91.29 90.99
TCP 90.57 91.37 90.97
Ours 90.10 91.57 90.83
$\Delta$0.47 +0.20 -0.14
(b) ImageNet
@@
(e) EuroSAT
Method Base Novel HM
CoOp 92.19 54.74 68.69
CoCoOp 87.49 60.04 71.21
TCP 91.63 74.73 82.32
Ours 95.70 82.37 88.54
$\Delta$ +4.07 +7.64 +6.21
(h) StanfordCars
Method Base Novel HM
CoOp 78.12 60.40 68.13
CoCoOp 70.49 73.59 72.01
TCP 80.80 74.13 77.32
Ours 81.97 76.10 78.93
$\Delta$ +1.17 +1.97 +1.60
(k) FGVCAircraft
Method Base Novel HM
CoOp 40.44 22.30 28.75
CoCoOp 33.41 23.71 27.74
TCP 41.97 34.43 37.83
Ours 45.67 36.50 40.57
$\Delta$ +3.70 +2.07 +2.75
(c) Caltech101
@@
CoOp 84.69 56.05 67.46
CoCoOp 82.33 73.45 77.67
TCP 87.13 80.77 83.83
Ours 87.13 83.10 85.07
$\Delta$ +0.00 +2.33 +1.24
(i) Flowers102
@@
(l) SUN397
@@
TABLE II: Comparison between InPK and the existing state-of-the-art methods in the base-to-novel generalization setting. Our method achieves the best accuracy in both base and novel classes, demonstrating the strong generalization ability of our method.
CoOp [4] CoCoOp [5] PromptSRC [7] HPT [12] ProMetaR [44] TCP  [41] CoPrompt [14] MMA [30] Ours
(IJCV22) (CVPR22) (ICCV23) (AAAI24) (CVPR24) (CVPR24) (ICLR24) (CVPR24)
Base 82.69 80.47 84.26 84.32 84.39 84.13 84.00 83.20 85.17
Novel 63.22 71.69 76.10 76.86 76.93 75.36 77.23 76.80 78.79
HM 71.66 75.83 79.97 80.23 80.49 79.51 80.48 79.87 81.85
Figure 3: Comparison of InPK with existing methods in the few-shot learning setting. All models are trained with 1, 2, 4, 8, and 16 shots per class and deployed on the full test set. Our method shows competitive performance in few-shot learning, achieving the highest average results.
TABLE III: Comparison of InPK and existing methods in the cross-data evaluation setting. All models are trained with 16-shot samples on the source dataset ImageNet and evaluated on 10 other datasets. Our method achieves the highest average accuracy, indicating that InPK has excellent robustness to different data distributions.
@@
Few-shot classification. To evaluate the model’s ability to learn with extremely limited supervision, we assess its performance across few-shot scenarios. Specifically, the model is trained on a small number of labeled images (1, 2, 4, 8, and 16 shots per class) and evaluated on a dataset containing the same classes as those in the training samples.
@@
Domain generalization. To assess the model’s robustness on out-of-distribution datasets, we evaluate the model trained on ImageNet-1K across four ImageNet variants that exhibit different types of domain shifts: ImageNet-V2, ImageNet-Sketch, ImageNet-A, and ImageNet-R.
@@
[…]
IV-B Base-to-Novel Generalization
@@
Table I presents the performance of InPK across 11 recognition datasets in the base-to-novel generalization setting. We implement all methods using a few-shot training approach, with 16 randomly sampled shots per base class. Compared to the recently proposed method TCP [41], our approach shows enhanced performance on 10 out of 11 datasets in terms of harmonic mean (HM), with only a slight drop in performance on the Food101 dataset relative to method TCP. For the average result, since we infuse class-specific prior knowledge into learnable tokens in Pki, these tokens are expected to learn the common features shared across modalities. This allows our method to perform better on novel class, improving the average accuracy of novel class by 3.43%. Additionally, InPK maintains strong performance on the base class as it can infer more discriminative text features, with an increase of 1.04% in accuracy. When considering both base and novel classes, InPK achieves an absolute average gain of 2.34% over TCP in terms of HM, demonstrating an effective balance between in-domain and out-of-domain data performance.
@@
Table II presents a comparison of the average performance of our method against more recent prompt-tuning methods across 11 datasets. The results evident that our method outperforms other methods on both base and novel classes, which demonstrates that the InPK not only maintains excellent recognition accuracy on base class but also exhibits strong generalization capabilities for novel class. Moreover, HPT [12] and CoPrompt [14] are the recently proposed methods that introduce prior knowledge. It can be seen that the performance of these methods that introduce prior knowledge but are not related to learnable tokens is lower than that of our proposed method, which verifies the importance of infusing prior knowledge into learnable tokens.
IV-C Few-Shot Learning
We further investigate the robustness of our InPK in the few-shot learning setting. Figure 3 presents the performance of InPK across different shots (1, 2, 4, 8, and 16 shots per class) on 11 datasets. We compare InPK with recent prompt tuning methods, including CoOp, CoCoOp, and TCP, achieving the best results on the majority of datasets. Compared to the previous state-of-the-art method TCP, InPK achieves an average improvement of 1.74% across different shot settings. Notably, the performance gain becomes more pronounced as the number of shots decreases. This is attributed to the incorporation of class-specific prior knowledge, which serves as additional informative cues to compensate for the scarcity of labeled samples, thereby enhancing the model’s ability to perceive fine-grained differences. This demonstrates that InPK maintains high performance even in limited data scenarios.
IV-D Cross-Dataset Evaluation
@@
To further verify the generalization ability of the model, we evaluate the model and state-of-the-art methods in a cross-data setting. As shown in Table III, our model exhibits outstanding generalization on 8 out of the 10 datasets, with a significant improvement over other methods. The average accuracy reaches 69.90%, demonstrating strong robustness across varying data distributions. This is because the model infuses the prior knowledge into learnable tokens at various feature levels, which guides the model to focus on category information and better discover latent representations of novel class attributes, thereby improving the model’s performance on unseen classes.
TABLE IV: Comparison of InPK with existing methods in the domain generalization evaluation setting.
Source Target
ImageNet ImageNetV2 ImageNet-S ImageNet-A ImageNet-R Average
CoOp 71.51 49.71 75.21 47.99 64.20 59.28
CoCoOp 71.02 50.63 76.18 48.75 64.07 59.91
CoPrompt 70.80 50.50 77.51 49.43 64.25 60.42
MMA 71.00 51.12 77.32 49.13 64.33 60.48
TCP 71.20 51.20 76.73 49.50 64.60 60.51
Ours 71.77 51.07 77.47 49.27 64.63 60.61
IV-E Domain Generalization
To evaluate the model’s robustness on out-of-distribution datasets, we conduct a domain generalization evaluation. Table IV compares our method with existing approaches in this setting. InPK demonstrates competitive performance, achieving the highest average accuracy across the four ImageNet variant datasets.
IV-F Ablation Study
@@
Impact of prompt length. We study the effect of prompt length $M$ on model performance in a base-to-novel generalization setting. As shown in Figure 4 (left), changing $M$ has minimal impact on base class performance. However, for the novel class, performance improves as the prompt length increases. Ultimately, we select $M=6$, which yields the highest harmonic mean (HM) accuracy, as the optimal prompt length for the model.
@@
Impact of the number of attribute words. We analyze the impact of the number of attribute words $N$ on model performance in the base-to-novel generalization setting as shown in Figure 4 (right). By varying the number of attribute words, we observe the accuracy of the base class, new class, and harmonic mean (HM). Overall, as the number of attribute words increases, the HM index gradually improves. This is because richer class-specific textual knowledge provides more comprehensive supervision for the model. Notably, when the number of attribute words $N=25$, the HM value is the highest, indicating that this amount of detailed prior knowledge optimizes the balance between base and novel class performance.
@@
Impact of inserting attribute-aware attention into different Transformer layers. Table V presents the analysis of inserting attribute-aware attention into different Transformer layers within PKi in the base-to-new generalization setting. The results indicate that even with $J=1$, where attribute-aware attention is only applied in the initialization stage, the model’s performance shows a significant improvement. This highlights the importance of providing a prior knowledge-based initialization for the learnable tokens. The model performance is observed to improve with the increase of the prompt depth of module insertion, which can be attributed to the fact that class-specific prior knowledge is able to reduce the loss of category information. Additionally, inserting attribute-aware attention beyond the ninth layer decreases performance, likely because earlier layers have already captured high-level features. Further insertions may interfere with these representations, making the model overly sensitive and reducing its generalization ability. The optimal result is obtained by incorporating attribute-aware attention into the transformer layers up to a depth $J=9$.
TABLE V: Impact of insert layers.
@@
[…]
TABLE VI: Comparison of different prompt initialization strategies in the context of base-to-novel generalization, with average results across 11 datasets.
Initialization Base Novel HM
[…]
Impact of prior knowledge infusion strategy. We experimented with strategies for infusing prior knowledge into learnable tokens, including ‘Concat’, ‘MLP’, and ‘AAA’. In the ‘Concat’ strategy, The prior knowledge and the learnable tokens are unrelated, they are concatenated and sent to the subsequent encoder layers. In the ‘MLP’ strategy, the two are passed through a fully connected layer to produce updated learnable tokens. Lastly, in the ‘AAA’ strategy, they are processed through the attribute-aware attention module to generate the new learnable tokens. Experimental results in Table VII show that the ‘AAA’ strategy achieves the best performance. This is because ‘Concat’ or’ MLP’ tends to incorporate irrelevant information directly, leading to a decline in the quality of the fused representation. In contrast, the attribute-aware attention module employs a weight distribution strategy to automatically filter out irrelevant or unimportant details, reducing the influence of redundancy and noise. This enables the model to selectively focus on the most valuable aspects of prior knowledge, resulting in fused learnable tokens that are both more discriminative and generalizable.
TABLE VII: Comparison of different prior knowledge infusion strategies for base-to-novel generalization across 11 datasets.
Infusion Strategy Base Novel HM
Concat 83.63 77.95 80.69
MLP 85.02 78.23 81.48
AAA 85.17 78.79 81.85
Analysis of model components. We conduct a comprehensive analysis of the key components of InPK, as illustrated in Table VIII. Starting from the baseline CoOp [4] (#1), which fine-tunes VLMs with learnable tokens, we observe poor performance on novel classes due to overfitting to base classes. When prior knowledge is introduced to the model (#2), we observe an improvement of 7.2% of HM. This gain is attributed to the richer class-specific information provided by prior knowledge, offering more comprehensive supervision. Due to the overlap of common attribute words among different classes, class confusion often occurs. To address this, regularization $L_{text}$ is introduced (#3), further improving the model’s performance by 1.08%. Building on this, we integrate the attribute-aware attention module (#4), which infuses prior knowledge into the learnable tokens and facilitates their interaction with prior knowledge across multiple feature levels. This enhancement allows the model to explicitly focus on class-relevant information, resulting in a performance boost of 0.86%. Finally, we propose a learnable text-to-vision projection layer to accommodate the adjustments in the text branch (#5), specifically the introduction of prior knowledge. These additions ensure better visual and textual semantics alignment by simultaneously refining both branches. As a result, the model achieves superior performance on both base and novel classes, with the HM improving by 10.19% compared to the baseline. In conclusion, the contribution of each model component to performance improvement is highlighted.
TABLE VIII: Further analysis of the model components in the base-to-novel generalization setting across 11 datasets. Results are shown for base classes (B), novel classes (N), and harmonic mean (HM).
Dataset #1 #2 #3 #4 #5
B 76.47 76.50 76.17 77.30 77.60
ImNet N 67.88 70.40 72.37 71.73 71.37
HM 71.92 73.32 74.22 74.41 74.35
B 98.00 98.00 98.33 98.33 98.67
Caltech N 89.81 94.43 95.13 95.33 94.30
HM 93.73 96.18 96.70 96.81 96.44
B 93.67 94.40 95.07 94.47 94.67
Pets N 95.29 94.57 97.50 97.63 97.63
HM 94.47 94.57 97.50 97.63 97.63
B 78.12 78.07 77.50 80.57 81.97
Cars N 60.40 73.00 74.67 75.17 76.10
HM 68.13 75.45 76.06 77.78 78.93
B 97.60 97.53 97.50 98.43 98.17
Flowers N 59.67 77.80 82.50 80.27 81.27
HM 74.06 86.55 89.38 88.43 88.92
B 88.33 89.50 90.63 90.07 90.10
Food N 82.26 81.53 81.00 80.67 81.13
HM 85.19 89.73 90.91 90.81 90.83
B 40.44 40.17 38.83 43.50 45.67
Aircraft N 22.30 35.40 38.63 37.40 36.50
HM 28.75 37.63 38.73 40.22 40.57
B 80.60 80.60 80.27 82.57 82.77
SUN397 N 65.89 81.53 81.00 80.67 81.13
HM 72.51 81.06 80.63 81.61 81.94
B 79.44 81.10 80.87 82.60 84.43
DTD N 41.18 67.07 69.30 69.83 71.30
HM 54.24 73.42 74.64 75.68 77.31
B 92.19 91.10 90.47 91.63 95.70
EuroSAT N 54.74 63.23 64.93 71.77 82.37
HM 68.69 74.65 75.60 80.49 88.54
B 84.69 85.23 85.03 86.37 87.13
UCF N 56.05 79.50 82.93 83.30 83.10
HM 67.46 82.27 83.97 84.81 85.07
B 82.69 82.93 82.79 84.17 85.17
Average N 63.22 75.17 77.29 77.70 78.79
HM 71.66 78.86 79.94 80.80 81.85
@@
To further verify the contributions of both prior knowledge and attribute-aware attention to model performance, we conduct a comprehensive analysis of the key components in the text branch, as illustrated in Table IX. In this analysis, the learnable projection layer from text prompts to visual prompts is consistently retained. In this analysis, Know. represents the input attribute words, Attn. refers to the attribute-aware attention. In the text branch, when only class-specific prior knowledge is applied, the model exhibits improvement with an average gain from 79.03% to 80.69% due to the additional supervision provided by the rich attribute information. On the other hand, when only attribute-aware attention is used, the model strikes a better balance between base and novel classes with an average improvement of 0.92%. When both components (Know. and Attn.) are incorporated together, the model achieves substantial gains across both base and new classes (83.70% vs 85.17%, 74.86% vs 78.79%). These results demonstrate that model performance is not solely influenced by prior knowledge itself but also by the manner in which the prior knowledge is processed. Specifically, the combination of prior knowledge with learnable tokens through the attribute-aware attention mechanism plays a crucial role in enhancing the model’s performance.
TABLE IX: Ablation on attribute words, attribute-aware attention and learnable text-to-vision projection layer in InPK.
Text encoder Base Novel HM
Know. Attn.
@@
Effectiveness of prior knowledge infusion. To further assess the impact of prior knowledge infusion (pkin) into learnable tokens, we apply pkin only before the first encoder layer in method CoOp [4], KgCoOp [29], and PromptSRC [7]. As shown in Table X, all three methods exhibit significant improvement, confirming the effectiveness of pkin.
TABLE X: Comparison of baseline methods with and without pkin in the base-to-novel generalization setting. ‘pkin’ denotes prior knowledge infusion into learnable tokens before feeding them into the encoder layer.
@@
+ pkin 85.15+0.89 78.60+2.50 81.74+1.77
V Conclusion